Calculating the Predetermined Overhead Rate, Applying Overhead to Production At the beginning of the year, Debion...

  1. Calculating the Predetermined Overhead Rate, Applying Overhead to Production

    At the beginning of the year, Debion Company estimated the following:

    Overhead $881,200
    Direct labor hours 93,000

    Debion uses normal costing and applies overhead on the basis of direct labor hours. For the month of March, direct labor hours were 19,530.

    Required:

    1. Calculate the predetermined overhead rate for Debion. Round your answer to the nearest cent.
    $ per direct labor hour

    2. Calculate the overhead applied to production in March. Round your intermediate calculations to two decimal places and round your answer to nearest whole value.
    $
Answer #1
1) Predetermined overhead
predetermined overhead = overhead/direct labor hour
predetermined overhead = 881200/93000 = 9 per hour (rounded off)
2) overhead applied = direct labor hour*predetermine overhead
overhead applied = 19530*9 = 175770

Please give a thumbs up if it is helpful & let me know if any doubt